# SLOPE and Designing Robust Studies for Generalization

### v1

@@ 4.2 Estimation
Our development of SLOPE has been based on the sensitivity model (3). While our reasons of choosing this model have been discussed in Section 2.2, we further discuss potential extensions of SLOPE to other sensitivity models.
To begin with, we remark on extending the exponential tilting shift, i.e., the right hand side of equation (3), to more general forms. As shown in Remark 7, this extension is straightforward; SLOPE can be still well defined and the connection between SLOPE and IF still holds in a more general form.
Remark 2 (The Direction of Sensitivity).
@@
We consider a broader class of sensitivity models that tilts the density ratio by a given non-negative function $\rho(O,X,\gamma)$ and incorporates the exponential tilting model (3) as a special case with $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$:
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}$ $\propto\rho(O,X,\gamma),$
where $\rho(O,X,0)=1$ and $\int\rho(O,X,\gamma)dP_{O\mid X}\in(0,\infty)$. Then the result in Theorem 3 becomes:
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{O,X}^{0}}\left(\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]\right),$
where $\dot{\rho}$ is the derivative of $\rho$ with respect to $\gamma$. Here, $\dot{\rho}$ characterizes the belief on how violations to conditional exchangeability may change the conditional distribution $Q_{O\mid X}^{\gamma}$. For example, when $\rho(O,X,\gamma)=\exp(\gamma\cdot O)$ as in (3), the derivative is $\dot{\rho}(O,X,0)=O$. Therefore, the term $\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\}\right]$ (i.e., $O-\mu(X)$ in (10)) can be viewed as the direction of sensitivity that’s unexplained by the observed variable $X$.
Next, we note that SLOPE under model (3) is based on the shift in a conditional density ratio. One important future direction is to generalize the concept of SLOPE, i.e., a derivative-based robustness measure that quantifies design-level robustness, to other types of sensitivity models. For example, in contrast to specifying $Q_{O\mid X}/P_{O\mid X}$ (semi-)parametrically, one may bound the difference between $Q_{O\mid X}$ and $P_{O\mid X}$ by some specified degree, which elicits a set identification instead of a point identification. Such extension is non-trivial due to the difficulty in (a) generalizing the notion of derivative from a point to a set, and (b) using the derivative as a metric to guide applications. We briefly illustrate the challenges in Remark 3 and defer details to Section J.5 of the Supplement.
@@
To fix ideas, consider Zeng et al., (2023)’s sensitivity model for $\psi^{\textnormal{mean}}$, which assumes the target conditional distribution, denoted as $Q_{O\mid X}^{\textnormal{bias}}$, deviate from $P_{O\mid X}$ by at most $\gamma$ in conditional means:
@@
Under (68), Zeng et al., (2023) showed that the target estimand is sharply bounded below and above by $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$ and $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$, respectively. Then, an analogous definition of SLOPE where we take the derivative of the upper and lower bounds with respect to $\gamma$ yields 1-1 and $1$, respectively. Since these two numbers disagree, the two-sided limit in Definition 1 no longer exists and the corresponding derivative is not well-defined. Moreover, even if we take the maximum magnitude of derivatives for the target estimand in (68), the resulting value (i.e., $1$) cannot be naturally interpreted by intrinsic properties of the design.
@@ 5.1 Data Background
We illustrate the use of SLOPE through a re-analysis of Banerjee et al., (2015)’s multi-national experiment that assesses the Graduation program in six countries (Ethiopia, Ghana, Honduras, India, Pakistan, and Peru). The Graduation program provides a holistic set of services, including asset transfers, consumption support and other career and health services, to poor households, with the purpose of improving their livelihoods. Between 2007 and 2014, in each village, eligible households were randomized to the intervention or control group by local non-profit organizations. The experiment lasted for 24 months, and data were collected before and after the program.
We adopt the potential outcome notation stated in Example 4 where $Y(a)$ denotes the potential outcome under treatment $a$, with $a=1$ denoting participation in the program and $a=0$ denoting otherwise. In subsections below, we use SLOPE to study the violation of conditional exchangeability in transporting the average potential outcome under treatment, i.e., $O=Y(1)$, from one country (i.e., the source population $P$) to another country (i.e., the target population $Q$). Since $O=Y(1)$ is a potential outcome in contrast to an observed variable, identification of the SLOPE requires additional causal assumptions, in specific stable unit treatment variable assumption (SUTVA) and strong ignorability, on the source population. See Section D in the Supplement for these assumptions and Section F.1 for their validity in this dataset.
We focus on two types of outcome variables, the per capita consumption and the physical health index, respectively, in Sections 5.2 and 5.3. For each outcome, the baseline covariate corresponds to the same variable measured prior to intervention. Our analyses are based on the complete data with overlapped baseline measurement (i.e., Assumption 1 holds). To harmonize the baseline covariate across countries and for simplicity, we discretize the measurement into coarse categorical variables as $X$ with consistent definitions. Note that due to the randomization of $A$, transformations of $X$ will not affect the plausibility of strong ignorability of $A$. For estimation, we have assumed that the conditional expectation of the outcome is linear in the baseline $X$ and the village where the household locates. Moreover, for SLOPE for the median in Section 5.2, the residual of the linear model is assumed to be normal. For details on data-preprocessing, estimation, and auxiliary results for Sections 5.2 and 5.3, see Sections F.3 and F.4, respectively, in the Supplement.
@@ 5.2 Which Source Country Is More Robust?
In this section, we study the SLOPE of transporting the per capita consumption across countries. The outcome variable is the per capita consumption averaged over measurements taken at two time points: immediately after the program and one year later. We adopted the logarithm transformation and after transformation, the outcome ranges from 1.4 to 6.4. We analyze complete data (83% of the sample) with overlapped baseline measurement (99% of the complete data) in five out of six countries. One country (Ethiopia) was excluded from this analysis because their consumption support, as part of the services during intervention, is substantially different from other countries. Sample sizes range from 771 to 2359 households across countries, with an average treatment rate of $35\%$. For each pair of countries, we treat one as the target and estimate the SLOPE of transporting the mean and median (i.e., $\psi^{\rm{\textnormal{mean}}}$ and $\psi^{{\textnormal{med}}}$) from the other (i.e., source).
Results are shown in Table 1. Given a target country and an estimand, the SLOPE is primarily determined by the data distribution in the source country. In most cases, using India or Peru as the source yields a lower SLOPE (i.e., a lower sensitivity) compared with using other source countries while using Pakistan yields a higher SLOPE (i.e., a higher sensitivity). This pattern of SLOPE remains consistent across estimands, i.e., source countries with a lower SLOPE of the mean usually also has a lower SLOPE of the median (e.g., India and Peru) and vice versa (e.g., Ghana). Overall, SLOPE suggests that among the six countries, India and Peru are more robust and Pakistan is least robust to generalize from. To understand the ranking in robustness across source countries, we recall from Sections 3.2 and 3.3 that the SLOPE for mean and median depends on the heterogeneity/dispersion of the source distribution. As shown by the box plot of $Y(1)$ in Figure 3, source countries with a higher SLOPE often have a uniformly higher spread of $Y(1)$ conditional on each category of $X$. Conversely, source countries with a lower SLOPE, for example India and Peru, have more concentrated values of $Y(1)$. A deeper contextual explanation for the robustness of India and Peru is beyond the scope of this work; we present one hypothesis in Section F.3 of the Supplement based on our understanding of the design of Banerjee et al., (2015).
Table 1: The estimated SLOPEs for transporting the counterfactual log-transformed per capita consumption under treatment from a source country (by rows) to a target country (by columns). Bootstrap standard errors are in the parentheses. For each target country and each estimand, the lowest SLOPE values among source countries are bold faced.
@@
In addition to comparing robustness across source countries, SLOPEs in Table 1 also present the comparison in robustness between the mean and the median. SLOPEs for mean and median are close and no one is dominating the other. In most cases the median has a smaller SLOPE, with exceptions where the its SLOPE is higher than the mean’s by no greater than 0.01. In addition, Section F.3.2 in the Supplement shows that the median and mean themselves (not their SLOPEs) have comparable estimated variances. Therefore, generally, we recommend using the median in transporting the average of per capita consumption.
Finally, since the outcome data are measured in the target country, we can assess the bias introduced by conditional exchangeability (i.e., “oracle bias”) empirically. This bias can also be approximated with SLOPE and an estimate of $\gamma$, in light of the first-order approximation in (5). Although bias approximation is not the designed/recommended usage of SLOPE, for completeness, we check the bias approximation with SLOPE by comparing it with the oracle bias. As shown by Figure 7 of the Supplement, the approximation works reasonably well. Additional details are in Section F.3.3 of the Supplement.
@@ 5.3 Which Health Index Is More Robust for Generalization?
Banerjee et al., (2015) constructed a physical health index to capture the variation in overall physical health of individuals in a household. Specifically, the index is an (equally weighted) average of three standardized variables (z-scores): did not miss work due to illness ($Y_{\text{notMiss}}(1)$), activities of daily living score ($Y_{\text{act}}(1)$), and perception of health status ($Y_{\text{perc}}(1)$). In this section, we ask whether there is another way to define a physical health index using SLOPE so that it’s less sensitive for generalization.
Formally, suppose we rewrite the physical health index as a weighted average of three z-scores:
$\alpha_{\text{notMiss}}Y_{\text{notMiss}}(1)+\alpha_{\text{act}}Y_{\text{act}}(1)+\alpha_{\text{perc}}Y_{\text{perc}}(1).$
@@
We focus on households in three countries (India, Ethiopia, Peru) where all three variables were measured prior to and after the treatment assignment. Our analysis is based on complete data (84% of all samples) with overlapped baseline measurement (80% of the complete data), which results in 730 households in India, 785 in Ethiopia, and 1307 in Peru with an average treatment rate of 43%.
Figure 4 presents the SLOPE across different weight combinations where the target country is Ethiopia and the source country is (left) India or (right) Peru. When the source country is India, the SLOPE is minimized at $\alpha_{\text{notMiss}}=0.10$, $\alpha_{\text{act}}=0.55$, $\alpha_{\text{perc}}=0.35$. Upon closer examination, the variable notMiss has the highest variance compared to the other two variables and thus, putting a low weight on it minimizes the overall variance of $O$. Also, when the source country is Peru, the SLOPE is minimized with the weight representing the left bottom vertex where $\alpha_{\text{notMiss}}=1$. Again, upon closer examination, this was because the variance of $Z_{\text{notMiss}}$ in Peru is substantially lower than the other two variables. More broadly, similar to the theoretical discussion from Section 3.2 and Corollary 1, SLOPE prefers distributions of the physical index that are more homogeneous and less variable in order to be less sensitive to generalization. While this phenomenon is intuitive, some readers may be concerned that generalizing from a homogeneous distribution may not be “useful”. We will address such concerns and further clarify the role of SLOPE in Section 6.
Figure 4: SLOPE for the mean of the health summary across weighting schemes. In each panel, edges of the triangle represent the weight in percentage scale of three z-scores. The contours represent the SLOPE where a lighter color means a lower SLOPE, which is more favorable. The star represents the case where the SLOPE is lowest in the simplex.
@@ 6 Discussion
This work proposes a measure, Sensitivity to LOcal Perturbation from Exchangeability (SLOPE), to quantify the change in the target estimand with respect to local violations of the conditional exchangeability assumption in transportability and generalizability. It reflects several design-level quantities of a generalization problem and can thus guide the design for generalization, including choices of source/target distributions and the target estimand. In particular, when interpreted with care, notably with respect to units (see Section 3), SLOPE can provide a simple and intuitive metric for choosing robust study designs for generalization under violations of conditional exchangeability. With SLOPE, under the validity of Assumption 1, some observations are useful in choosing robust study designs in practice:
@@
    With a fixed target population, a source population that’s more homogeneous in $P_{O\mid X}$ is more robust to violation of the conditional exchangeability. For instance, the mean’s SLOPE becomes smaller when $\sigma^{2}(X)$ reduces (Corollary 1), which is in agreement to Tipton and Olsen, (2018). As another example, in Section 5.2, countries with a smaller SLOPE (e.g., India and Peru as shown in Table 1) also have a less variant distribution of $O$ given $X$ (Figure 3).
@@
    With a fixed source population, a target population that contains more “homogeneous units” with respect to $P_{O\mid X}$ is more robust to generalize to. For example, in the left panel of Figure 2, SLOPE increases as the target population incorporates a higher proportion of units from a subgroup (defined with $X$) that corresponds to a more diverse $O$.
@@
    When holding populations fixed, a target functional is more robust when its influence function projects more (in proportion) onto the space of shared variable $X$. In a trivial example, as implied by Theorem 3, a target functional that concerns $X$ only, i.e., $\psi(Q_{O,X})=\psi(Q_{X})$, has zero SLOPE. Additionally, as shown in Section 5.3, for a weighted average of several physical health variables, increasing weights to variables that are better explained by $X$ will also reduce the magnitude of SLOPE, thereby improving robustness.
We re-emphasize that SLOPE is developed under the validity of the overlap assumption (i.e., Assumption 1), which means that the source population is already sufficiently large compared with the target population, in terms of $P_{X}$ and $Q_{X}$. In addition, while SLOPE is proposed under situations when conditional exchangeability is unlikely to hold, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) on the importance of guaranteeing conditional exchangeability through domain expertise and careful data collection (e.g., by collecting a rich $X$). However, when it is unrealistic to meet the conditional exchangeability with the observed set of $X$, SLOPE is a useful tool to assess sensitivity/robustness of the underlying design, and to guide designs in future generalization problems.
Finally, we highlight some extensions and future directions. First, while this paper focuses on the conditional exchangeability assumption in generalizations, by defining $P$ and $Q$ differently, SLOPE naturally extends to measuring the sensitivity for other types of conditional exchangeability assumptions in causal inference and missing data; see Section J.4 of the Supplement. Second, an interesting example in Section J.6 of the Supplement connects SLOPE mathematically to the marginal interventional effect proposed by Zhou and Opacic, (2022) with the incremental propensity score intervention (Kennedy,, 2019), highlighting the potential analogies of SLOPE in incremental treatment effects. Lastly, while our result has been based on sensitivity model (3), the high level idea of SLOPE as a derivative-based robustness measure for study designs is “generalizable”. As stated in Section 4.3, an important direction is to extend SLOPE from the specific model (3) to other sensitivity models that posit less parametric assumptions.

### v2

@@ 4.3 Extending SLOPE to Other Types of Sensitivity Analysis
Our development of SLOPE is based on the sensitivity model (3). While our reasons of choosing this model have been discussed in Section 2.2, we discuss two potential extensions of SLOPE to other sensitivity models.
First, consider an extension of (3) where we replace the exponential tilting term with a more general, non-negative function $\rho(O,X,\gamma)$. As we show below, SLOPE can still be well-defined with respect to $\rho$ and it maintains its analytic connection to the IF.
Theorem 4 (SLOPE and IF for $\rho$-Based Sensitivity Model).
@@
Consider a broader class of sensitivity models that tilts the density ratio by a non-negative function $\rho(O,X,\gamma)$:
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}$ $\propto\rho(O,X,\gamma),$ (14)
where $\rho(O,X,0)=1$ and $\int\rho(O,X,\gamma)OdP_{O\mid X}<\infty$. Suppose $\rho$ is differentiable at $\gamma=0$ with its derivative denoted as $\dot{\rho}(O,X,0)$ and Condition 7 in the Supplement holds. Then
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{X}}\left\{\textnormal{E}_{P_{O\mid X}}\left(\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X\}\right]\right)\right\}.$
Compared to Theorem 3, the relationship between IF and SLOPE defined under the sensitivity model (14) is driven by $\dot{\rho}$ at the point where conditional exchangeability holds (i.e., $\gamma=0$). Specifically, the residual variation in $\dot{\rho}$, as measured by $\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X\}$ can be viewed as the subspace defined in the sensitivity analysis that is not explained by $X$. When $\rho(O,X,\gamma)=\exp(\gamma O)$ as in our original sensitivity model (3), the derivative at $\gamma=0$ is $\dot{\rho}(O,X,0)=O$ and the subspace is defined by $O-\mu(X)$.
Second, an important future direction is to generalize SLOPE to “bound-based” sensitivity analysis (e.g., Rosenbaum, (1987); Tan, (2006); Zeng et al., (2023)) where the upper bound on the difference between $Q_{O\mid X}$ and $P_{O\mid X}$ is some function of $\gamma$. However, we believe such an extension is non-trivial due to the difficulty in (a) generalizing the notion of derivative from a point to a set, and (b) the most natural generalization of this set-based derivative may not be as insightful for guiding robust designs compared to our current approach. We briefly illustrate points (a) and (b) in Remark 2 and defer details to Section J.4 of the Supplement.
@@
To fix ideas, consider Zeng et al., (2023)’s sensitivity analysis of $\psi^{\textnormal{mean}}$ in the target population. Their sensitivity analysis assumes the target conditional distribution, denoted as $Q_{O\mid X}^{\textnormal{bias}}$, deviate from $P_{O\mid X}$ by at most $\gamma$ where the deviation is measured in terms of conditional means:
@@
Under (15), Zeng et al., (2023) showed that the target estimand $\psi^{\textnormal{mean}}$ is sharply bounded below and above by $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$ and $\gamma+\psi^{{\textnormal{mean}}}(Q_{O,X}^{0})$, respectively. Then, an analogous definition of SLOPE where we take the derivative of the upper and lower bounds with respect to $\gamma$ yields 1-1 and $1$, respectively. Since these two numbers disagree, the two-sided limit in Definition 1 no longer exists and the corresponding derivative is not well-defined.
Also, even if we take the maximum magnitudes of the two derivatives to resolve the issue (i.e., the “worst-case” SLOPE), we believe the resulting value (i.e., $1$) cannot be meaningfully interpreted as an intrinsic property of the study design because any source distribution $P_{O\mid X}$ or target distribution $Q_{X}$ will yield the same maximum of $1$. As mentioned in Section 3.2, a measure of robustness that is constant irrespective of the source or the target distribution does not align with some empirical recommendations on robust study designs for generalization. In contrast, our SLOPE based on either exponential tilting (3) or its generalized form (14) depends on $P_{O\mid X}$ and $Q_{X}$, and different choices of these distributions reflect differences in robustness between study designs.
@@ 5.1 Data Background
We illustrate how to use SLOPE to inform robust study designs by re-analyzing Banerjee et al., (2015)’s multi-national experiment. The goal of the experiment was to evaluate the Graduation program in six countries (Ethiopia, Ghana, Honduras, India, Pakistan, and Peru). The Graduation program provides a holistic set of services, including asset transfers, consumption support and other career and health services, to poor households. Between 2007 and 2014, eligible households in each village were randomized to the intervention or the control group and the experiment lasted for 24 months.
We adopt the potential outcome notation stated in Example 4 where $Y(a)$ denotes the potential outcome under treatment $a$, with $a=1$ denoting participation in the program and $a=0$ denoting otherwise. In subsections below, we use SLOPE to study the violation of conditional exchangeability in transporting the potential outcome of treatment, i.e., $O=Y(1)$, from one country (i.e., the source population $P$) to another country (i.e., the target population $Q$). We remark that identification of the SLOPE requires additional causal assumptions in the source population (i.e., stable unit treatment variable assumption [SUTVA] and strong ignorability), on the source population and these assumptions are satisfied because a randomized experiment was conducted; see Sections D and F.1 in the Supplement for details.
We focus on two types of outcome variables, the per capita consumption and the physical health index, respectively, in Sections 5.2 and 5.3. For each outcome, the baseline covariate corresponds to the same variable measured prior to intervention. Our analyses are based on complete data with overlapped baseline measurement (i.e., Assumption 1 holds). To harmonize $X$ across countries for generalizability, we discretize $X$ into coarse categorical variables; note that due to the randomization of $A$, these transformations of $X$ will not affect the plausibility of strong ignorability of $A$. Finally, for estimation, we have assumed that the conditional expectation of the outcome is linear in the baseline $X$ and the village where the household locates. Moreover, for SLOPE for the median in Section 5.2, the residual of the linear model is assumed to be normal for simplicity and diagnostics in Section F.3 of the Supplement suggest that this is a reasonable assumption. see Sections F.3 and F.4 in the Supplement.
@@ 5.2 Which Source Country Is More Robust?
In this section, we study the SLOPE of transporting the per capita consumption across countries. The outcome variable is the log of the average of per capita consumption at two time points and ranges from 1.4 to 6.4. One country (Ethiopia) was excluded from the analysis because their consumption support was substantially different from other countries. For each pair of countries, we treat one as the target population, the other as the source population, and estimate the SLOPE of the mean and median (i.e., $\psi^{\rm{\textnormal{mean}}}$ and $\psi^{{\textnormal{med}}}$).
Results are shown in Table 1. Given a target country, the SLOPE is primarily determined by the data distribution in the source country and does not vary much between the mean and the median. Using India or Peru as the source population yields a lower SLOPE (i.e., a lower sensitivity) compared with using other source countries. Also, there is minimal difference in the SLOPEs of the mean and median, although the median’s SLOPEs are consistently, but slightly, lower than the mean’s SLOPEs. Section F.3.2 in the Supplement further shows that the estimated median and mean themselves (not their corresponding SLOPEs) have comparable estimated variances and thus, there is no (empirical) loss in efficiency when we estimate the median. Therefore, from this analysis based on SLOPE, we generally recommend using the median in transporting the average of per capita consumption.
To explain why India and Peru generally have smaller SLOPEs, Figure 3 plots the the distribution of $Y(1)$ given some discrete values of $X$ across countries. We see that Ghana, which is a source country that almost always has the highest SLOPE, has higher spread of $Y(1)$ conditional on each category of $X$. Conversely, India and Peru, which are the source countries with lower SLOPEs, have more concentrated values of $Y(1)$ relative to others.
We conducted some sanity checks of our recommendation and we provide a summary of them. First, because the outcome data is actually measured in the target country, we are in a unique situation where we can assess whether the sensitivity model (3) is reasonable for our analysis. Specifically, we can assess the bias from violating conditional exchangeability empirically and check whether SLOPE can approximate this bias using the first-order approximation in (5). Figure 7 in Section F.3.3 of the Supplement shows that the first-order approximation works well where the bias can be well-approximated by SLOPE. Second, a deeper subject-matter expertise explanation for why India and Peru are robust compared to other countries with respect to violation of conditional exchangeability is beyond the scope of this work. Nevertheless, we present one hypothesis in Section F.3 of the Supplement based on our understanding of Banerjee et al., (2015).
Table 1: The estimated SLOPEs for transporting the counterfactual log-transformed per capita consumption under treatment (i.e., $O=Y(1)$) from a source country (by rows) to a target country (by columns). Bootstrap standard errors are in the parentheses. For each target country and estimand, the lowest SLOPEs among the source countries are bold faced.
@@ 5.3 Which Health Index Is More Robust for Generalization?
Banerjee et al., (2015) constructed a physical health index to capture the overall physical health of individuals in a household. Specifically, the index is an (equally weighted) average of three standardized variables (z-scores): did not miss work due to illness ($Y_{\text{notMiss}}(1)$), activities of daily living score ($Y_{\text{act}}(1)$), and perception of health status ($Y_{\text{perc}}(1)$). In this section, we ask whether there is another way to define the physical health index so that it’s less sensitive for generalization. Formally, suppose we rewrite the physical health index as a weighted average of three z-scores:
$O=\alpha_{\text{notMiss}}Y_{\text{notMiss}}(1)+\alpha_{\text{act}}Y_{\text{act}}(1)+\alpha_{\text{perc}}Y_{\text{perc}}(1).$
@@
We focus on households in three countries (India, Ethiopia, Peru) where all three variables that make up the index were measured. Also, like before, we filtered households so that the overlap was plausible.
Figure 4 presents the SLOPE across different weight combinations where the target country is Ethiopia and the source country is (left) India or (right) Peru. When the source country is India, the SLOPE is minimized at $\alpha_{\text{notMiss}}=0.10$, $\alpha_{\text{act}}=0.55$, $\alpha_{\text{perc}}=0.35$. Upon closer examination, the variable notMiss has the highest variance compared to the other two variables and thus, putting a low weight on it minimizes the overall variance of $O$. Also, when the source country is Peru, the SLOPE is minimized with the weight representing the left bottom vertex where $\alpha_{\text{notMiss}}=1$. Again, upon closer examination, this was because the variance of $Z_{\text{notMiss}}$ in Peru is substantially lower than the other two variables. More broadly, similar to the theoretical discussion from Section 3.2 and Corollary 1, SLOPE prefers distributions of the physical index that are less variable in order to be less sensitive to generalization.
Figure 4: SLOPE for the mean of the physical health index across different weights. In each panel, edges of the triangle represent the weight in percentage scale of the three z-scores that make up the index. The contours represent the SLOPE where a lighter color means a lower SLOPE. The star represents the point in the simplex where SLOPE is minimized.
@@ 6 Discussion
This main contribution of this paper is a simple and novel measure, SLOPE, to design robust studies for generalization. Specifically, SLOPE is inspired by principles from robust statistics and is a derivative-based metric that measures the change in the target estimand when there is a local violation of the conditional exchangeability assumption. SLOPE depends on two design-level quantities, (a) the target estimand and (b) the source and target distributions $P_{O\mid X}$ and $Q_{X}$, respectively. Changing either of these quantities can result in a different SLOPE and thus, researchers can assess the robustness of competing study designs for generalization.
Inspired by Tipton and Olsen, (2018)’s recommendations for robust generalization, we summarize some advice for designing robust studies based on SLOPE. We remark that these principles are assuming that Assumption 1 holds:
@@
    If the target population is fixed and the target estimand is the mean of $O$, a source population will be less sensitive to local violations of conditional exchangeability if the spread of $O$ given $X$ is small. For instance, the mean’s SLOPE becomes smaller when $\sigma^{2}(X)$ becomes smaller (Corollary 1). From our data example in Section 5.2, countries with a smaller SLOPE (e.g., India and Peru as shown in Table 1) also have a less variable distribution of $O$ given $X$ (Figure 3).
@@
    If the source population is fixed and the target estimand is the mean of $O$, it’s essential to understand which region of $X$ leads to the least amount of variation in $O$ in the source population. Once this region of $X$ is identified, a target population will be robust for generalization if it is homogeneous with respect to its $X$ and its $X$s focus around this region. More concretely, Section 3.2 and Corollary 1 discuss an example where $Q_{X}$ is selected to focus on regions with the least variability in $O$.
@@
    If both the source and the target populations are fixed, it’s less sensitive to choose a target estimand whose influence function projects more (in proportion) onto the space of shared variable $X$. In a trivial example inspired by Theorem 3, a target functional that concerns $X$ only, i.e., $\psi(Q_{O,X})=\psi(Q_{X})$, has zero SLOPE. Additionally, as shown in Section 5.3, for a weighted average of several physical health variables, increasing weights to variables that are better explained by $X$ will also reduce the magnitude of SLOPE, thereby improving robustness.
We re-emphasize that SLOPE is developed under the validity of the overlap assumption (i.e., Assumption 1), which means that the source population is already sufficiently large compared with the target population, in terms of $P_{X}$ and $Q_{X}$. In addition, we echo Tipton and Olsen, (2018) and Degtiar and Rose, (2023) on the general importance of guaranteeing conditional exchangeability through discussion with domain experts and careful data collection processes (e.g., by collecting a rich $X$). However, when it is infeasible or impractical to plausibly satisfy conditional exchangeability with the observed set of $X$, we believe SLOPE is a useful tool to assess the sensitivity/robustness of the underlying study design, and to guide future designs for generalization.
Finally, we highlight some extensions and future directions. First, while this paper focuses on violations of conditional exchangeability in generalizability, by defining $P$ and $Q$ differently, SLOPE naturally extends to measuring the sensitivity of conditional exchangeability in causal inference and missing data problems; see Section J.3 of the Supplement for these extensions. Second, Section J.5 of the Supplement connects SLOPE to the marginal interventional effect proposed by Zhou and Opacic, (2022) with the incremental propensity score intervention (Kennedy,, 2019), highlighting some potential mathematical connections between SLOPE and incremental treatment effects. Third, while our result has been based on sensitivity model (3), the high level idea of SLOPE as a derivative-based robustness measure for study designs is “generalizable.” As stated in Section 4.3, an important direction is to extend SLOPE to bound-based sensitivity models where a properly defined SLOPE not only exists, but also provides useful insights about robust study designs.